OpenStudy (anonymous):

4/(x-5)+1/(x+2)=2x+7/(x-5)(x+2)

OpenStudy (saifoo.khan):

shift the whole equation in 1 side. then take the LCM = (x-5)(x+2) solve.

OpenStudy (anonymous):

5x + 3 = 2x + 7 3x = 4 x=4/3